ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

9. Ordinary Differential Equations and Dynamical Systems

1991 MS Classification: 30 Schleicher, Dierk, Technische Universität München On iterated exponential maps Holomorphic dynamics has been a success story with the last 20 years. Best understood is the iteration theory for polynomials, in particular for those with a single critical point. We will discuss attempts to carry theory and results over to the family of iterated exponential maps z ↦ λe^{z}. This yields a description of parameter space similar to Mandelbrot and Multibrot sets, but with significant differences.